OpenStudy (anonymous):

If f(x) = −3x − 3 and g(x) = 2x + 1, find f/g(−2).

OpenStudy (jhannybean):

if we have g(x) = 2x +1 g(-2) can be found by plugging in -2 into the x-values for the g(x) function. g(-2) = 2(-2)+1 = ?

OpenStudy (anonymous):

−3 3 −1 1

OpenStudy (anonymous):

those are the answer choices. i got -1?

OpenStudy (jhannybean):

what does 2(-2)+1 =?

OpenStudy (anonymous):

-3

OpenStudy (jhannybean):

Good. now we're solving for f/(-3) So \[\large \frac{f(x)}{-3} = \frac{-3x-3}{-3}\]\[\large \frac{-3x}{-3}-\frac{3}{3}\] can you evaluate this now? :)

OpenStudy (anonymous):

x-1?

OpenStudy (jhannybean):

Good job!

OpenStudy (jhannybean):

\[\large \frac{f(x)}{g(-2)}= x-1\]

OpenStudy (anonymous):

so is the answer -1?

OpenStudy (jhannybean):

Nope, the answer is what you stated, x-1.

OpenStudy (jhannybean):

I don't believe you're evaluating the function for x, are you? If you are,in that case you would solve for x. x-1 = 0 add 1 to both sides x -1 +1 = 1 + 0 x = 1